OpenStudy (anonymous):

Nathan Alexander wants to purchase ears of corn for a family reunion picnic. A farmers’ market near his home offers six ears of corn for $1.25, 12 ears of corn for $2.64, or 36 ears of corn for $7.20. Which is the best buy?

OpenStudy (anonymous):

This is a very easy problem. Divide the cost by the number of ears of corn to figure out the unit price of each ear. See the results below. $1.25/6=0.21 per ear. $2.64/12=0.22 per ear $7.20/36=0.20 Therefore, the 36 ears of corn at $7.20 is the best buy

OpenStudy (anonymous):

A 16-inch cheese pizza at Fine Dough Pizza costs $10.00 and each additional topping is $1.50. The same size pizza at Johnson's Pizza costs $8.00 and each additional topping is $2.00. After how many toppings do the pizzas cost the same? Which pizza costs more with additional toppings?

OpenStudy (anonymous):

After 4 toppings. The Johnsons pizza costs more with additional toppings.